How many samples of size n=3 can be selected from a population with the following sizes: N=4, N=8, N=20, N=50? A population consists of the five numbers 2, 3, 6, 8 and 11. Consider samples of size 2 that can be drawn from this population.


1
Expert's answer
2022-04-11T16:11:45-0400

The number of possible samples which can be selected without replacement is

"\begin{pmatrix}
 N \\
 n
\end{pmatrix}=\cfrac{N! } {n! \cdot(N-n)! }."



"\begin{pmatrix}
 4\\
 3
\end{pmatrix}=\cfrac{4! } {3! \cdot1!} =4."



"\begin{pmatrix}
 8\\
 3
\end{pmatrix}=\cfrac{8! } {3! \cdot5! }=\cfrac{6\cdot7\cdot8}{2\cdot3}=56 ."



"\begin{pmatrix}
 20\\
 3
\end{pmatrix}=\cfrac{20! } {3! \cdot17! }=\cfrac{18\cdot19\cdot20}{2\cdot3}=1140 ."



"\begin{pmatrix}
 50\\
 3
\end{pmatrix}=\cfrac{50! } {3! \cdot47! }=\cfrac{48\cdot49\cdot50}{2\cdot3}=19600 ."



"\begin{pmatrix}
 5\\
 2
\end{pmatrix}=\cfrac{5! } {2! \cdot3! }=\cfrac{4\cdot5}{2}=10."

The possible samples:

(2, 3), (2, 6), (2, 8), (2, 11), (3, 6),

(3, 8), (3, 11), (6, 8), (6, 11), (8, 11).